# Bayesian Optimization of Robustness Measures Using Randomized GP-UCB-based Algorithms under Input Uncertainty

### v1

@@ 2 Preliminary
Let $f:\mathcal{X}\times\Omega\to\mathbb{R}$ be an expensive-to-evaluate black-box function, where $\mathcal{X}$ and $\Omega$ are finite sets [^1] We discuss the case where $\mathcal{X}$ and $\Omega$ are continuous sets in Appendix A.. For each $({\bm{x}}_{t},{\bm{w}}_{t})\in\mathcal{X}\times\Omega$, $f({\bm{x}}_{t},{\bm{w}}_{t})$ is observed with noise $\varepsilon_{t}$ as $y_{t}=f({\bm{x}}_{t},{\bm{w}}_{t})+\varepsilon_{t}$, where $\varepsilon_{1},\ldots,\varepsilon_{t}$ are mutually independent and follow some distribution with mean 0. Let ${\bm{x}}\in\mathcal{X}$ be a design variable and ${\bm{w}}\in\Omega$ be an environmental variable, where ${\bm{w}}$ is an uncontrollable variable and follows a distribution $P^{\ast}$. In black-box function optimization including environmental variables, two types of settings for environmental variables exist: simulator settings (Cakmak et al.,, 2020; Nguyen et al., 2021b, ; Beland and Nair,, 2017; Iwazaki et al., 2021a, ) and uncontrollable settings (Kirschner et al.,, 2020; Inatsu et al., 2024a, ; Inatsu et al.,, 2022; Iwazaki et al., 2021b, ). In the simulator setting, we can arbitrarily determine the value of ${\bm{w}}$ during optimization, whereas in the uncontrollable setting, we cannot control the value of ${\bm{w}}$ even during optimization. In the main body, we consider the simulator setting, and the uncontrollable setting is discussed in Appendix B. For a function $\vartheta({\bm{w}})$ that can be expressed as a function of ${\bm{w}}$, let $\rho(\cdot)$ be a function that maps $\vartheta({\bm{w}})$ to real numbers. In particular, when ${\bm{x}}$ is fixed, we define $\rho(f({\bm{x}},{\bm{w}}))\equiv F({\bm{x}})$ as a robustness measure of $f$ in ${\bm{x}}$. For example, representative robustness measures include the expectation measure $F({\bm{x}})=\mathbb{E}[f({\bm{x}},{\bm{w}})]$, worst-case measure $F({\bm{x}})=\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})$, and value-at-risk measure $F({\bm{x}})=\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}(f({\bm{x}},{\bm{w}})\leq b)\}$, where the expectation or probability is taken with respect to ${\bm{w}}$. In this paper, our goal is to identify the following ${\bm{x}}^{\ast}$ with as few function evaluations as possible:
@@
We emphasize that while the optimization target is $F({\bm{x}})$, we cannot observe $F({\bm{x}})$, but can only observe $f({\bm{x}},{\bm{w}})$ which contains noise.
@@
We introduce regularity assumptions for $f$. Let $k:(\mathcal{X},\Omega)\times(\mathcal{X},\Omega)\to\mathbb{R}$ be a positive-definite kernel satisfying $k(({\bm{x}},{\bm{w}}),({\bm{x}},{\bm{w}}))\leq 1$ for any $({\bm{x}},{\bm{w}})\in\mathcal{X}\times\Omega$. Assume that $f$ is a sample path from a GP $\mathcal{G}\mathcal{P}(0,k(({\bm{x}},{\bm{w}}),({\bm{x}}^{\prime},{\bm{w}}^{\prime})))$ with mean function 0 and kernel $k(\cdot,\cdot)$. Also assume that $\varepsilon_{t}$ follows the normal distribution with mean 0 and variance $\sigma^{2}_{{\rm noise}}$, where $f,\varepsilon_{1},\ldots,\varepsilon_{t}$ are mutually independent.
@@
In this study, we predict $F$ based on a surrogate model of the black-box function $f$. We assume that the prior distribution of $f$ is the GP $\mathcal{G}\mathcal{P}(0,k(({\bm{x}},{\bm{w}}),({\bm{x}}^{\prime},{\bm{w}}^{\prime})))$. In this case, given a dataset $\{({\bm{x}}_{j},{\bm{w}}_{j},y_{j})\}_{j=1}^{t}$, the posterior distribution of $f$ is again a GP, and its posterior mean $\mu_{t}({\bm{x}},{\bm{w}})$ and posterior variance $\sigma^{2}_{t}({\bm{x}},{\bm{w}})$ can be calculated from the properties of GPs (Williams and Rasmussen,, 2006) as follows:
@@
where ${\bm{k}}_{t}({\bm{x}},{\bm{w}})$ is the $t$-dimensional vector whose $j$-th element is $k(({\bm{x}},{\bm{w}}),({\bm{x}}_{j},{\bm{w}}_{j}))$, ${\bm{y}}_{t}=(y_{1},\ldots,y_{t})^{\top}$, ${\bm{I}}_{t}$ is the $t\times t$ identity matrix, and ${\bm{K}}_{t}$ is the $t\times t$ matrix whose $(j,k)$-th element is $k(({\bm{x}}_{j},{\bm{w}}_{j}),({\bm{x}}_{k},{\bm{w}}_{k}))$.
@@ 3 Proposed Method
In this section, we propose a BO method to efficiently identify ${\bm{x}}^{\ast}$. First, in Section 3.1, we construct credible intervals for $F({\bm{x}})$ based on credible intervals for $f({\bm{x}},{\bm{w}})$. Next, in Section 3.2, we give a method for estimating the optimal solution. In Section 3.3, we give a method for selecting ${\bm{x}}_{t}$ and ${\bm{w}}_{t}$. The pseudo-code of the proposed algorithm is given in Algorithm 1.
Algorithm 1 RRGP-UCB for robustness measures
@@
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}_{t}$ by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (3.4)
   Select the next evaluation point $\bm{w}_{t}$ by (3.5)
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
@@
3.1 Credible Interval for Robustness Measure
For each $({\bm{x}},{\bm{w}})\in\mathcal{X}\times\Omega$ and $t\geq 1$, let $Q_{t-1}({\bm{x}},{\bm{w}})=[l_{t-1}({\bm{x}},{\bm{w}}),u_{t-1}({\bm{x}},{\bm{w}})]$ be a credible interval for $f({\bm{x}},{\bm{w}})$, where $l_{t-1}({\bm{x}},{\bm{w}})$ and $u_{t-1}({\bm{x}},{\bm{w}})$ are give by
@@ 3.1 Credible Interval for Robustness Measure
Here, $\beta_{t}\geq 0$ is a user-specified trade-off parameter. Due to the properties of GPs, $f({\bm{x}},{\bm{w}})$ after data observation follows the normal distribution with mean $\mu_{t-1}({\bm{x}},{\bm{w}})$ and variance $\sigma^{2}_{t-1}({\bm{x}},{\bm{w}})$. Therefore, by setting $\beta_{t}$ appropriately  [^2] For example, if $\beta^{1/2}_{t}=1.96$, then $f({\bm{x}},{\bm{w}})\in Q_{t-1}({\bm{x}},{\bm{w}})$ holds with probability $0.95$., $Q_{t-1}({\bm{x}},{\bm{w}})$ contains $f({\bm{x}},{\bm{w}})$ with high probability. Next, for each ${\bm{x}}$, we define $G_{t-1}({\bm{x}})$, the set of functions with respect to ${\bm{w}}$, as follows:
$G_{t-1}({\bm{x}})=\{g({\bm{x}},{\bm{w}})\mid^{\forall}{\bm{w}}\in\Omega,g({\bm{x}},{\bm{w}})\in Q_{t-1}({\bm{x}},{\bm{w}})\}.$
@@
If $Q_{t-1}({\bm{x}},{\bm{w}})$ is a high-probability credible interval of $f({\bm{x}},{\bm{w}})$ for any ${\bm{w}}\in\Omega$, then the function $f({\bm{x}},{\bm{w}})$ on ${\bm{w}}$ is included in $G_{t-1}({\bm{x}})$ with high probability. Therefore, the following inequality holds with high probability:
$\inf_{g({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}}))\leq\rho(f({\bm{x}},{\bm{w}}))=F({\bm{x}})\leq\sup_{g({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}})).$ (3.1)
@@
We can construct a high-probability credible interval for $F({\bm{x}})$ by using the left and right sides of (3.1). However, we must emphasize that exact calculations of the left and right sides of (3.1) are not easy in general. For this reason, we introduce the lower bound ${\rm lcb}_{t-1}({\bm{x}})$ and upper bound ${\rm ucb}_{t-1}({\bm{x}})$ satisfying
${\rm lcb}_{t-1}({\bm{x}})\leq\inf_{g({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}})),\ \sup_{g({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}}))\leq{\rm ucb}_{t-1}({\bm{x}}).$ (3.2)
Inatsu et al., 2024a showed that ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ for commonly used robustness measures, including the expectation measure, can be analytically calculated using $l_{t-1}({\bm{x}},{\bm{w}})$ and $u_{t-1}({\bm{x}},{\bm{w}})$. We give representative robustness measures and their corresponding ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ in Table 3. Using this we define the credible interval $Q_{t-1}({\bm{x}})=[{\rm lcb}_{t-1}({\bm{x}}),{\rm ucb}_{t-1}({\bm{x}})]$ for $F({\bm{x}})$.
3.2 Estimation of the Optimal Solution

### v2

@@ 2 Preliminary
Let $f:\mathcal{X}\times\Omega\to\mathbb{R}$ be an expensive-to-evaluate black-box function, where $\mathcal{X}$ and $\Omega$ are finite sets [^1] The case where $\mathcal{X}$ and $\Omega$ are continuous is discussed in Appendix A.. For each $({\bm{x}}_{t},{\bm{w}}_{t})\in\mathcal{X}\times\Omega$, we observe $f({\bm{x}}_{t},{\bm{w}}_{t})$ with noise $\varepsilon_{t}$ as follows: $y_{t}=f({\bm{x}}_{t},{\bm{w}}_{t})+\varepsilon_{t}$, where $\varepsilon_{1},\ldots,\varepsilon_{t}$ are mutually independent and follow some distribution with zero mean. Let ${\bm{x}}\in\mathcal{X}$ be a design variable and ${\bm{w}}\in\Omega$ an environmental variable, where ${\bm{w}}$ is uncontrollable and follows a distribution $P^{\ast}$. In black-box optimization involving environmental variables, two types of settings are considered: simulator settings (Cakmak et al.,, 2020; Nguyen et al., 2021b, ; Beland and Nair,, 2017; Iwazaki et al., 2021a, ) and uncontrollable settings (Kirschner et al.,, 2020; Inatsu et al., 2024a, ; Inatsu et al.,, 2022; Iwazaki et al., 2021b, ). In the simulator setting, the value of ${\bm{w}}$ can be arbitrarily selected during optimization, whereas in the uncontrollable setting, ${\bm{w}}$ cannot be controlled even during optimization. In the main text, we focus on the simulator setting; the uncontrollable setting is discussed in Appendix B. Let $\vartheta({\bm{w}})$ be a function of ${\bm{w}}$, and let $\rho(\cdot)$ be a mapping from the function $\vartheta(\cdot)$ to the real numbers. For any function $\varphi({\bm{x}},{\bm{w}})$ defined on $\mathcal{X}\times\Omega$, when ${\bm{x}}$ is fixed, $\varphi({\bm{x}},{\bm{w}})$ is a function of ${\bm{w}}$ and is denoted by $\varphi({\bm{x}},\cdot)$. In particular, when ${\bm{x}}$ is fixed, we define $\rho(f({\bm{x}},\cdot))\equiv F({\bm{x}})$ as a robustness measure of $f$ in ${\bm{x}}$. Representative robustness measures include: the expectation measure $F_{1}({\bm{x}})=\mathbb{E}[f({\bm{x}},{\bm{w}})]$, the worst-case measure $F_{2}({\bm{x}})=\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})$, the best-case measure $F_{3}({\bm{x}})=\sup_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})$, the $\alpha$-value-at-risk measure $F_{4}({\bm{x}};\alpha)=\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}(f({\bm{x}},{\bm{w}})\leq b)\}$, the $\alpha$-conditional value-at-risk measure $F_{5}({\bm{x}};\alpha)=\mathbb{E}[f({\bm{x}},{\bm{w}})|f({\bm{x}},{\bm{w}})\leq F_{4}({\bm{x}};\alpha)]$, and the mean absolute deviation measure $F_{6}({\bm{x}})=\mathbb{E}[|f({\bm{x}},{\bm{w}})-F_{1}({\bm{x}})|]$, where the expectation or probability is taken with respect to ${\bm{w}}$. Our goal is to identify the following ${\bm{x}}^{\ast}$ using as few function evaluations as possible:
@@
We emphasize that while the optimization target is $F({\bm{x}})$, we cannot directly observe $F({\bm{x}})$; instead, only the noisy evaluations of $f({\bm{x}},{\bm{w}})$ are available.
@@
We introduce regularity assumptions for the function $f$. Let $k:(\mathcal{X},\Omega)\times(\mathcal{X},\Omega)\to\mathbb{R}$ be a positive-definite kernel such that $k(({\bm{x}},{\bm{w}}),({\bm{x}},{\bm{w}}))\leq 1$ for all $({\bm{x}},{\bm{w}})\in\mathcal{X}\times\Omega$. Assume that $f$ is a sample path from a GP $\mathcal{G}\mathcal{P}(0,k(({\bm{x}},{\bm{w}}),({\bm{x}}^{\prime},{\bm{w}}^{\prime})))$ with zero mean and kernel function $k(\cdot,\cdot)$. We further assume that the noise terms $\varepsilon_{t}$ are independently drawn from a normal distribution with mean zero and variance $\sigma^{2}_{{\rm noise}}$, and that $f,\varepsilon_{1},\ldots,\varepsilon_{t}$ are mutually independent.
@@
In this study, we predict $F$ based on a surrogate model of the black-box function $f$. We assume that the prior distribution of $f$ is a GP $\mathcal{G}\mathcal{P}(0,k(({\bm{x}},{\bm{w}}),({\bm{x}}^{\prime},{\bm{w}}^{\prime})))$. Given the dataset $\{({\bm{x}}_{j},{\bm{w}}_{j},y_{j})\}_{j=1}^{t}$, the posterior distribution of $f$ remains a GP. The posterior mean $\mu_{t}({\bm{x}},{\bm{w}})$ and posterior variance $\sigma^{2}_{t}({\bm{x}},{\bm{w}})$ are given by standard results from the GP regression (Williams and Rasmussen,, 2006):
@@
where ${\bm{k}}_{t}({\bm{x}},{\bm{w}})$ is the $t$-dimensional vector whose $j$-th element is $k(({\bm{x}},{\bm{w}}),({\bm{x}}_{j},{\bm{w}}_{j}))$, ${\bm{y}}_{t}=(y_{1},\ldots,y_{t})^{\top}$, ${\bm{I}}_{t}$ is the $t\times t$ identity matrix, and ${\bm{K}}_{t}$ is the $t\times t$ kernel matrix with the $(j,k)$-th element $k(({\bm{x}}_{j},{\bm{w}}_{j}),({\bm{x}}_{k},{\bm{w}}_{k}))$.
Notations
We summarize particularly important notations used in the main text in Table 3.
Table 3: Notations used in the main text and their meanings.
Notation Meaning
${\bm{x}}$ Design variable
${\bm{w}}$ Environmental variable
$\mathcal{X}$ Set of design variables
$\Omega$ Set of environmental variables
$f({\bm{x}},{\bm{w}})$ Black-box function defined on $\mathcal{X}\times\Omega$
$\varepsilon_{t}$ Observetion noise following a normal distribution
$\sigma^{2}_{{\rm noise}}$ Variance of the noise ditribution
$\varphi({\bm{x}},{\bm{w}})$ Arbitrary function defined on $\mathcal{X}\times\Omega$
$\varphi({\bm{x}},\cdot)$ In $\varphi({\bm{x}},{\bm{w}})$, a function with respect to ${\bm{w}}$ when ${\bm{x}}$ is fixed
$\rho(\cdot)$ Robustness measure, a mapping from functions with respect to ${\bm{w}}$ to real numbers
$F({\bm{x}})$ Target robustness measure defined by $F({\bm{x}})=\rho(f({\bm{x}},\cdot))$
${\bm{x}}^{\ast}$ Optimal solution defined by $\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}F({\bm{x}})$
$k(\cdot,\cdot)$ Kernel function defined on $(\mathcal{X}\times\Omega)\times(\mathcal{X}\times\Omega)$
$\mu_{t}({\bm{x}},{\bm{w}})$ Posterior mean based on GP for $f({\bm{x}},{\bm{w}})$
$\sigma^{2}_{t}({\bm{x}},{\bm{w}})$ Posterior variance based on GP for $f({\bm{x}},{\bm{w}})$
$u_{t}({\bm{x}},{\bm{w}})$ Upper confidence bound for $f({\bm{x}},{\bm{w}})$
$l_{t}({\bm{x}},{\bm{w}})$ Lower confidence bound for $f({\bm{x}},{\bm{w}})$
$\beta_{t}$ Parameter for adjusting the confidence bound width
${\rm ucb}_{t}({\bm{x}})$ Upper confidence bound for $F({\bm{x}})$
${\rm lcb}_{t}({\bm{x}})$ Lower confidence bound for $F({\bm{x}})$
$\hat{\bm{x}}_{t}$ Estimated solution for the optimal solution defined by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},\cdot))$
$(\cdot)_{+}$ Operator defined by $(a)_{+}=\max\{a,0\}$
$\tilde{\bm{x}}_{t}$ Optimistic maximum solution defined by $\tilde{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}({\rm ucb}_{t-1}({\bm{x}})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}}))_{+}$
${\bm{x}}_{t}$ Next design variable to be evaluated
$\xi_{t}$ Random variable following the chi-squared distribution with two digrees of freedom
${\bm{w}}_{t}$ Next environmental variable to be evaluated
$r_{t}$ Instantaneous regret regarding the optimal solution and estimated solution
$R_{t}$ Cumulative instantaneous regret regarding the optimal solution and estimated solution
$\gamma_{t}$ Maximum information gain
$\mathbb{E}_{t-1}[\cdot]$ Conditional expectation given the dataset $D_{t-1}=\{({\bm{x}}_{1},{\bm{w}}_{1},\varepsilon_{1},\beta_{1}),\ldots,({\bm{x}}_{t-1},{\bm{w}}_{t-1},\varepsilon_{t-1},\beta_{t-1})\}$
$\hat{t}$ Optimal index given by $\hat{t}=\operatornamewithlimits{argmax}_{1\leq i\leq t}\mathbb{E}_{t-1}[F(\hat{\bm{x}}_{i})]$
@@ 3 Proposed Method
In this section, we propose a BO method to efficiently identify ${\bm{x}}^{\ast}$. First, in Section 3.1, we construct credible intervals for $F({\bm{x}})$ based on credible intervals for $f({\bm{x}},{\bm{w}})$. Next, in Section 3.2, we present a method for estimating the optimal solution. In Section 3.3, we describe a method for selecting ${\bm{x}}_{t}$ and ${\bm{w}}_{t}$. The pseudo-code of the proposed algorithm is provided in Algorithm 1.
Algorithm 1 RRGP-UCB for robustness measures.
@@
   Generate $\xi_{t}$ from chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}_{t}$ by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho({\mu_{t-1}({\bm{x}},\cdot)})$
   Select next evaluation point $\bm{x}_{t}$ by (3.4)
   Select next evaluation point $\bm{w}_{t}$ by (3.5)
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding observed data
@@
3.1 Credible Interval for Robustness Measures
For each $({\bm{x}},{\bm{w}})\in\mathcal{X}\times\Omega$ and $t\geq 1$, let $Q_{t-1}({\bm{x}},{\bm{w}})=[l_{t-1}({\bm{x}},{\bm{w}}),u_{t-1}({\bm{x}},{\bm{w}})]$ denote a credible interval for $f({\bm{x}},{\bm{w}})$, where $l_{t-1}({\bm{x}},{\bm{w}})$ and $u_{t-1}({\bm{x}},{\bm{w}})$ are given by
@@ 3.1 Credible Interval for Robustness Measures
Here, $\beta_{t}\geq 0$ is a user-defined trade-off parameter. Due to the properties of GPs, the posterior distribution of $f({\bm{x}},{\bm{w}})$ after observing data is a normal distribution with mean $\mu_{t-1}({\bm{x}},{\bm{w}})$ and variance $\sigma^{2}_{t-1}({\bm{x}},{\bm{w}})$. Therefore, by selecting an appropriate value of $\beta_{t}$  [^2] For example, if $\beta^{1/2}_{t}=1.96$, then $f({\bm{x}},{\bm{w}})\in Q_{t-1}({\bm{x}},{\bm{w}})$ holds with probability $0.95$., the interval $Q_{t-1}({\bm{x}},{\bm{w}})$ contains $f({\bm{x}},{\bm{w}})$ with high probability. Next, for each ${\bm{x}}$, we define $G_{t-1}({\bm{x}})$, the set of functions over ${\bm{w}}$, as follows:
$G_{t-1}({\bm{x}})=\{{g({\bm{x}},\cdot)}\mid{\rm for\ any}\ {\bm{w}}\in\Omega,g({\bm{x}},{\bm{w}})\in Q_{t-1}({\bm{x}},{\bm{w}})\}.$
@@
If $Q_{t-1}({\bm{x}},{\bm{w}})$ is a high-probability credible interval for $f({\bm{x}},{\bm{w}})$ for all ${\bm{w}}\in\Omega$, then the function $f({\bm{x}},\cdot)$ lies in $G_{t-1}({\bm{x}})$ with high probability. Therefore, the following inequality holds with high probability:
$\inf_{{g({\bm{x}},\cdot)}\in G_{t-1}({\bm{x}})}\rho({g({\bm{x}},\cdot)})\leq\rho({f({\bm{x}},\cdot)})=F({\bm{x}})\leq\sup_{{g({\bm{x}},\cdot)}\in G_{t-1}({\bm{x}})}\rho({g({\bm{x}},\cdot)}).$ (3.1)
@@
We can thus construct a high-probability credible interval for $F({\bm{x}})$ using the left- and right-hand sides of (3.1). However, computing the exact bounds in (3.1) is generally intractable. To address this, we introduce the lower bound ${\rm lcb}_{t-1}({\bm{x}})$ and upper bound ${\rm ucb}_{t-1}({\bm{x}})$, which satisfy:
${\rm lcb}_{t-1}({\bm{x}})\leq\inf_{{g({\bm{x}},\cdot)}\in G_{t-1}({\bm{x}})}\rho({g({\bm{x}},\cdot)}),\ \sup_{{g({\bm{x}},\cdot)}\in G_{t-1}({\bm{x}})}\rho({g({\bm{x}},\cdot)})\leq{\rm ucb}_{t-1}({\bm{x}}).$ (3.2)
Inatsu et al., 2024a showed that, for commonly used robustness measures, including expectation, the bounds ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ can be analytically calculated using $l_{t-1}({\bm{x}},{\bm{w}})$ and $u_{t-1}({\bm{x}},{\bm{w}})$ [^3] In Tables 3 and 4 in Inatsu et al., 2024a , the terms “risk measure” and “Bayes risk” are used instead of “robustness measure” and “expectation measure,” respectively.. Inatsu et al., 2024a differs from this study in that it deals with multi-objective optimization for robustness measures under input uncertainty and adopts the assumption that black-box functions are in RKHS. On the other hand, it shares commonalities with this study in that it adopts a GP as a surrogate model and calculates upper and lower bounds for robustness measures based on (3.2). Therefore, the computational methods for ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ presented in Inatsu et al., 2024a for various robustness measures are applicable to the setting of this study. Representative robustness measures and their corresponding ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ are summarized in Table 4. On the other hand, even for robustness measures not listed in Table 4, it is possible to approximate ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ by sampling. First, we generate $S$ sample paths $f^{(1)}({\bm{x}},{\bm{w}}),\ldots,f^{(S)}({\bm{x}},{\bm{w}})$ from the posterior distribution of the function $f({\bm{x}},{\bm{w}})$ and then estimating ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ as follows:
${\rm lcb}_{t-1}({\bm{x}})=\min_{1\leq s\leq S}\rho(f^{(s)}({\bm{x}},\cdot)),\ {\rm ucb}_{t-1}({\bm{x}})=\max_{1\leq s\leq S}\rho(f^{(s)}({\bm{x}},\cdot)).$
Using these bounds, we define the credible interval $Q_{t-1}({\bm{x}})=[{\rm lcb}_{t-1}({\bm{x}}),{\rm ucb}_{t-1}({\bm{x}})]$ for $F({\bm{x}})$.
3.2 Estimation of Optimal Solution